OpenStudy (anonymous):

What value of x will make the equation 2x + 6x= 56 true?

OpenStudy (anonymous):

2x+6x=56 => x(2+6)=56 => x=56/8 => x=7 so for every x=7 the equation is true

OpenStudy (anonymous):

8x=56 x=7

OpenStudy (idily101):

8x=56 8 8 x= 7